solve by method of substitution
3x-6y=21
3x+14=y


Answer by kyawkhant(2)

Solution:
3x-6y= 21.........eqn(1)
3x+14= y.........eqn(2)
Substitute y in eqn (1)
3x-6(3x+14)=21
3x-18x-84=21
-15x=21+84
-15x=105
x=105/-15
x=-7
x=-7, y=3(-7)+14
y=-21+14
y= -7
Therefore, x=-7 and y=-7